OpenStudy (anonymous):

Help plz: Solve for the variable in the equation 17 = t + 10

OpenStudy (anonymous):

17=t+10 17-10=t+10-10 7=t t=7

OpenStudy (anonymous):

Thanks man

OpenStudy (anonymous):

No problem